A ball is thrown vertically upward with an intitial velocity of 100 feet per second.
a) Determine a function h(t) that gives the height of the ball at time t seconds after it is thrown. The height function is h(t)=-16t^2+vt+h where v is the initial velocity of the ball and h is the initial height of the ball.
b) how long does it take the ball to reach its maximum height?
c) what is the maximum height the ball reaches?

Answer by stanbon(48564)

A ball is thrown vertically upward with an initial velocity of 100 feet per second.
a) Determine a function h(t) that gives the height of the ball at time t seconds after it is thrown. The height function is h(t)=-16t^2+vt+h where v is the initial velocity of the ball and h is the initial height of the ball.
--------
If you assume the ball was thrown upward from the ground the equation is
Ans: h(t)=-16t^2+100t because h=0.
--------------------
b) how long does it take the ball to reach its maximum height?
Maximum occurs when t = -b/2a = -100/(2*-16) = 3.125 seconds.
---------------------
c) what is the maximum height the ball reaches?
h(3.125) = -16(3.125)^2+100(3.125) = 156.25 feet
======================
Cheers,
Stan H.